# Enhanced DeepONet for 1-D consolidation operator learning: an architectural investigation

### v1

@@ Abstract
Deep Operator Networks (DeepONets) have emerged as a powerful surrogate modeling framework for learning solution operators in PDE-governed systems. While their use is expanding across engineering disciplines, applications in geotechnical engineering remain limited. This study systematically evaluates several DeepONet architectures for the one-dimensional consolidation problem. We initially consider three architectures: a standard DeepONet with the coefficient of consolidation embedded in the branch net (Models 1 and 2), and a physics-inspired architecture with the coefficient embedded in the trunk net (Model 3). Results show that Model 3 outperforms the standard configurations (Models 1 and 2) but still has limitations when the target solution (excess pore pressures) exhibits significant variation. To overcome this limitation, we propose a Trunknet Fourier feature-enhanced DeepONet (Model 4) that addresses the identified limitations by capturing rapidly varying functions. All proposed architectures achieve speedups ranging from 1.5 to 100 times over traditional explicit and implicit solvers, with Model 4 being the most efficient. Larger computational savings are expected for more complex systems than the explored 1D case, which is promising. Overall, the study highlights the potential of DeepONets to enable efficient, generalizable surrogate modeling in geotechnical applications, advancing the integration of scientific machine learning in geotechnics, which is at an early stage.
@@
[label1] organization=School of Civil and Environmental Engineering, Atlanta, Georgia Institute of Technology, city=Atlanta, postcode=30332, state=GA, country=USA
@@ 1 Introduction
Recent advances in scientific machine learning (Faroughi et al., 2022) have introduced computationally efficient deep learning-based surrogate modeling frameworks (Huang et al., 2025) to approximate the solutions of PDEs obtained using numerical solvers. Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019; Cuomo et al., 2022) constitute one of these emerging frameworks. PINNs approximate PDE solutions by enforcing governing equations as soft constraints within the loss function during neural network training. Although PINNs have been widely explored in various engineering domains, their application in geotechnical engineering remains relatively limited, with recent studies focusing on assessing the potential of PINN. For example, Bekele (2021) evaluated the PINN potential for forward and inverse problems in soil consolidation. Wang et al. (2024) evaluated PINN’s performance in the context of experimental consolidation data. Gao et al. (2023); Yang et al. (2025) applied PINN to seepage analysis and soil-water characteristic curves. Liu et al. (2025) assessed PINN’s potential in seismic site response analysis, leveraging Fourier feature embedding.
@@
Although PINNs have shown promise for solving PDEs in many engineering disciplines, recent studies (Grossmann et al., 2024; Bueno et al., 2025) indicate that they are not necessarily more effective PDE solvers compared to traditional numerical methods. PINNs are designed to approximate solutions for a fixed set of PDE coefficients and boundary conditions, which means that they require retraining if those conditions change (Faroughi et al., 2022). This retraining process is computationally expensive (Sharma and Shankar, 2022), which poses challenges in the use of PINNs in applications with variable system configurations. Additionally, PINNs require an explicit formulation of governing equations, which limits their applicability in domains where physical behavior is described by empirical or phenomenological models rather than well-defined PDEs.
@@
Neural operators (Kovachki et al., 2023; Li et al., 2020a; Lu et al., 2021; Li et al., 2020b) have emerged as an alternative in scientific machine learning that addresses some of the PINN’s limitations. Neural operators are a new type of deep learning framework designed to learn operators that map between infinite-dimensional function spaces. Unlike conventional deep learning methods that approximate solutions for specific inputs, neural operators learn the underlying solution operator that maps input functions (e.g., material property fields, boundary conditions, loadings) to output solution fields. This operator-based learning enables neural operators to generalize across varying initial and boundary conditions without retraining. Additionally, neural operators do not require explicit knowledge of governing equations, which makes them particularly well-suited for complex physical phenomena where a closed-form PDE is not defined. Compared to conventional numerical methods or PINNs, which need to solve the underlying PDEs for each new set of conditions (Koric et al., 2024; Jiang et al., 2024; Faroughi et al., 2022), neural operators offer significantly more efficient inference based on learned function mapping. These advantages make them particularly well-suited for surrogate models in engineering applications with complex operating conditions.
@@
Among various neural operator architectures, Deep Operator Networks (DeepONet), proposed by Lu et al. (2021), have gained substantial attention due to their formulation. DeepONet is based on the universal approximation theorem for operators (Chen and Chen, 1995), which establishes its ability to approximate nonlinear solution operators of PDEs given sufficient network complexity and data coverage. This framework allows DeepONet to efficiently learn generalizable operators (Lu et al., 2022).
@@
DeepONet can achieve greater data efficiency, faster training convergence, and improved accuracy (Lu et al., 2021) compared to conventional architectures, such as feed-forward networks (FNNs), sequence-to-sequence models (seq2seq), and residual networks (ResNets) in operator learning tasks. Furthermore, the architecture of DeepONet offers interpretability through its explicit separation of functional roles: the branch net learns to encode the input function, representing the coefficients of a latent basis, while the trunk net learns to output corresponding basis functions over the spatio-temporal domain, which are then combined with these coefficients to form the solution. This design mirrors the structure of functional mappings in infinite-dimensional spaces, where solution operators can be viewed as weighted combinations of basis functions parameterized by input functions (Naylor and Sell, 1982). This formulation is in alignment with the mathematical framework of operator learning (Kovachki et al., 2023).
@@
DeepONet has been applied in various fields, such as fluid mechanics (Kumar et al., 2024; Bai et al., 2024), material science (Yin et al., 2022), structural mechanics (He et al., 2023; Ahmed et al., 2024), and thermodynamics (Koric and Abueidda, 2023; Kushwaha et al., 2024); these efforts have highlighted DeepONet’s potential in modeling complex physical processes. However, the potential of DeepONet in geotechnical engineering remains largely unexplored. This study aims to introduce DeepONet as a surrogate modeling approach for geotechnical problems, focusing on the one-dimensional (1D) consolidation phenomenon, one of the most prevalent geotechnical engineering problems. The motivation for exploring the consolidation problem is that it can serve as a benchmark to demonstrate DeepONet’s capabilities to the geotechnical community, while also highlighting the role of the different architectural adaptations explored in this study.
@@
More specifically, this study presents a systematic investigation into DeepONet architectural choices tailored for the consolidation phenomenon. We first explain the standard architecture of DeepONet, and then explore three model learning architectures to get initial insights. The first two models follow DeepONet variants investigated previously (Tan and Chen, 2022; Jin et al., 2022), and the third model adopts a non-traditional architectural choice, with its better accuracy supported by the analytical solution structure of the consolidation problem. Insights and challenges identified from the various explored architectures are shared and utilized to propose a DeepONet architecture enhanced with Trunknet Fourier feature embedding, which demonstrates superior performance. Overall, this study contributes to the integration of advanced scientific machine learning into geotechnical applications.
@@ 2 Deep operator networks(DeepONet)
DeepONet’s learned basis and corresponding coefficients eliminate the need for manually designed basis functions, which are typically required in traditional numerical methods. For example, the finite element method uses local basis functions defined over a mesh, while spectral methods rely on global functions like sine and cosine series (e.g., Fourier series) chosen in advance (Williams et al., 2024). This flexibility allows DeepONet to effectively capture complex, nonlinear operator mappings in high-dimensional PDEs.
@@ 3 Consolidation and associated PDEs
$u(z,t)=\sum_{m=0}^{\infty}\left[\frac{2u_{0}}{M}\sin\left(\frac{Mz}{H_{\mathrm{dr}}}\right)\right]e^{-M^{2}T_{v}}$ (5)
Where $M=\frac{\pi}{2}(2m+1)$ under single drainage boundary condition, $u_{0}$ is the initial excess pore pressure, $H_{dr}$ is the drainage path length, and $T_{v}=\frac{C_{v}t}{H_{dr}^{2}}$ is the dimensionless time factor. This analytical form reveals that the solution is composed of spatial sine functions modulated by exponential decay terms dependent on $C_{v}$ and time $t$. This insight informs one of the DeepONet architecture choices that is later explored.
@@ 4 Data generation and training
We use the finite difference method to numerically solve Equation 4. The solution obtained is used to train the DeepONet and to test the performance. The finite difference method first discretizes the spatial domain uniformly into $N_{z}$ nodes separated by a spacing $\Delta z$. It then approximates the spatial second derivative at each internal node $i$ ($21\leq i\leq N_{z}-$ 2), using a second-order central finite-difference scheme (Equation 6).
@@
When generating data, we consider both uniform and spatially varying initial excess PWP profiles $u(z,0)$ as the initial condition. For the uniform profile, we randomly sample constant functions between 10 kPa and 20 kPa. For the spatially varying (non-uniform) profiles, we sample functions from a Gaussian Random Field (GRF), denoted as $u(z,0)\sim\mathcal{G}(m,C(z_{1},z_{2}))$. The spatial correlation of this GRF is defined by a Gaussian covariance function:
$C(z_{1},z_{2})=\sigma^{2}\exp\left(-\frac{||z_{1}-z_{2}||^{2}}{l^{2}}\right)$ (9)
Here, $m$ is mean, $C(z_{1},z_{2})$ represents the covariance between two points at depths $z_{1}$ and $z_{2}$. The term $||z_{1}-z_{2}||$ is the Euclidean distance between these points. The parameter $l$ is the correlation length (or length-scale), which controls the smoothness of the sampled functions; a larger $l$ generates smoother profiles. The term $\sigma^{2}$ is the variance of the field, controlling the overall magnitude of the covariance. We use a $\sigma^{2}$ value of 1000 $kPa^{2}$, a correlation length of 0.5, and randomly select $m$ from 10 kPa to 20 kPa. The coefficient of consolidation, $C_{v}$, is sampled uniformly from $[0.3,1.0]$ m2/year. The spatial domain spans the normalized depth $\frac{z}{H_{dr}}\in[0,1]$ and normalized time $T_{v}=\frac{C_{v}t}{H_{dr}^{2}}\in[0,2]$.
@@
We train and validate DeepONet using N=40,000 and N=5,000 input functions, respectively. Each input function represents an initial excess PWP and is discretized at m=100 fixed, equally spaced sensor locations, denoted by $\{x_{\ell}\}_{\ell=1}^{100}$ (see the left side figure in Figure 2a). That is, for each input function $u^{(i)}$, the branch network receives the vector $\left[u^{(i)}(x_{1}),u^{(i)}(x_{2}),\dots,u^{(i)}(x_{100})\right]$ as input. The corresponding output functions (i.e., solutions $s=\mathcal{G}(u)$) are sampled at P=100 randomly selected evaluation points $y=(z,t)$ (see the red markers in the right side figure in Figure 2a), sampled independently for each input function. These locations are denoted $\{y_{j}^{(i)}\}_{j=1}^{100}$, resulting in a total of $40,000\times 100$ training examples of the form $\left(u^{(i)},y_{j}^{(i)},\mathcal{G}(u^{(i)})(y_{j}^{(i)})\right)$. The validation set is constructed in the same manner, with P=100 output evaluations per input, resulting in $5,000\times 100$ training examples. For testing, we assess the generalization capability of the learned operator $\mathcal{G}_{\Theta}$ on 500 unseen input functions during training. Each is evaluated over a dense $100\times 100$ uniformly spaced grid in the output domain (see the right side figure in Figure 2b), yielding P=10,000 output locations per function to visualize full solution fields. All input and output data are standardized using the mean and standard deviation computed from the training set.
@@ 5 Model architectures
    Model 1 (Figure 3): This model serves as our baseline and adheres to the standard DeepONet formulation introduced by Lu et al. (2021). The branch net processes the input function, which in this case consists of the discretized initial PWP profile, $u(z,0)$. The scalar coefficient of consolidation, $C_{v}$, is concatenated with the discretized $u(z,0)$ vector, forming a single input tensor for the branch net. This design philosophy treats $C_{v}$ as a global conditioning parameter. The implicit assumption is that the branch net should learn a combined representation of how the initial PWP distribution and the material property $C_{v}$ jointly influence the consolidation process. The trunk net takes the spatiotemporal coordinates $(z,t)$ as input.
@@
    Model 2 (Figure 4): This architecture modifies the branch net to handle $u(z,0)$ and $C_{v}$ separately before integrating their influence. One branch net is dedicated to encoding the initial PWP function, $u(z,0)$, while a second, auxiliary branch net processes the coefficient of consolidation, $C_{v}$, generating distinct latent representations for each. These two latent representations are then concatenated and passed through a ”merge net” (another small neural network) to produce the final branch output vector. This approach allows the model to potentially learn more disentangled features for the initial condition and the PDE coefficient before their combined effect on the solution basis is determined. The trunk net architecture remains the same, taking $(z,t)$ as input.
@@
The performance of Models 1, 2, and 3 is discussed in Section 6, highlighting the challenges to predict excess PWP distribution during the early stages of consolidation, especially for non-uniform distributions across depth. This motivated the authors to explore an additional option. Specifically, an additional architecture option considers Model 3 enhanced with Fourier Feature Embedding (Tancik et al., 2020), which can help to learn rapidly varying functions (e.g., Wang et al. (2021), Liu et al. (2025)).
@@
$\gamma(\textbf{v})=\begin{bmatrix}\sin{(2\pi\mathbf{B}\textbf{v}})\\ \cos{(2\pi\mathbf{B}\textbf{v}})\end{bmatrix}^{T}$ (11)
where v is the input feature vector (trunk net input in this problem) of length $k$ and $\gamma(\textbf{v})$ is the embedded Fourier feature vector with a length of $2m$. $\mathbf{B}$ is a $m$ by $k$ matrix of frequencies, sampled from $\mathcal{N}(\mathbf{0},\sigma^{2}I)$. The tunable hyperparameters $m$ and $\sigma$ are pre-specified before the neural network training. This embedding introduces periodic features that enrich the input representation, allowing neural networks to approximate functions with high variation. Leveraging the Fourier feature embedding, we propose the following architecture:
@@
All the networks (i.e., branch, trunk, and merge net) in DeepONet models in this study employ fully connected multilayer perceptrons (MLPs). Each MLP comprises 6 layers with 30 neurons per layer. The output dimension of the branch and trunk net for the dot product is set to $q=50$ ( Equation 3). For Fourier feature embedding hyperparameters (Equation 11) in model 4, we use $m=50$ and $\sigma=1.0$.

### v2

# Operator Learning for Consolidation: An Architectural Comparison for DeepONet Variants

@@ Abstract
Deep Operator Networks (DeepONets) have emerged as a powerful surrogate modeling framework for learning solution operators in PDE-governed systems. While their use is expanding across engineering disciplines, applications in geotechnical engineering remain limited. This study systematically evaluates several DeepONet architectures for the consolidation problem. We initially consider three architectures: a standard DeepONet with the coefficient of consolidation embedded in the branch net (Models 1 and 2), and a physics-inspired architecture with the coefficient embedded in the trunk net (Model 3). Results show that Model 3 outperforms the standard configurations (Models 1 and 2) but still has limitations when the target solution (excess pore pressures) exhibits significant variation. To overcome this limitation, we propose a Trunknet Fourier feature-enhanced DeepONet (Model 4) that addresses the identified limitations by capturing rapidly varying functions. We further extend Model 4 to 3D scenarios. Although the computational speedup can be modest in the 1D case (1.5-100× compared with traditional solvers), the speedup becomes more pronounced in 3D, reaching approximately 1,000×. Leveraging this efficiency, we offer a conceptual demonstration of DeepONet’s potential to accelerate uncertainty quantification in a 3D consolidation problem. Overall, the study highlights the potential of DeepONets to enable efficient, generalizable surrogate modeling in geotechnical applications, advancing the integration of scientific machine learning in geotechnics, which is at an early stage.
@@
[…]
@@ 1 Introduction
Recent advances in scientific machine learning (faroughi2022sciml) have introduced computationally efficient deep learning-based surrogate modeling frameworks (huang2025deeplearningpdes) to approximate PDE solutions obtained via numerical solvers. Physics-Informed Neural Networks (PINNs) (raissi2019pinn; cuomo2022sciml_pinn) constitute one of these emerging frameworks. PINNs approximate PDE solutions by enforcing governing equations as soft constraints within the loss function during neural network training. Although PINNs have been widely explored in various engineering domains, their application in geotechnical engineering remains relatively limited, with recent studies focusing on assessing the potential of PINN. For example, bekele2021pinn_consolidation_jrmge evaluated the PINN potential for forward and inverse problems in soil consolidation. gao2023pinn_seepage_phreatic_line; yang2025seepage_swcc applied PINN to seepage analysis and soil-water characteristic curves. liu2025leveraging assessed PINN’s potential in seismic site response analysis, leveraging Fourier feature embedding.
@@
Although PINNs have shown promise for solving PDEs across many engineering disciplines, recent studies (grossmann2024pinn_vs_fem; bueno2025pinn_vs_fdm) indicate that they are not necessarily more effective PDE solvers than traditional numerical methods. PINNs are designed to approximate solutions for a fixed set of PDE coefficients and boundary conditions, which means that they require retraining if those conditions change (faroughi2022sciml). This retraining process is computationally expensive (sharma2022accelerated_pinn), posing challenges for using PINNs in applications with variable system configurations. Additionally, PINNs require an explicit formulation of governing equations, which limits their applicability in domains where physical behavior is described by empirical or phenomenological models rather than well-defined PDEs.
@@
Neural operators (kovachki2023neural_operators; li2020fourierNOs; lu2021learning; li2020graphNOs) have emerged as an alternative in scientific machine learning that address some of PINNs’ limitations. Neural operators are a new type of deep learning framework designed to learn operators that map between infinite-dimensional function spaces. Unlike conventional deep learning methods that approximate solutions for specific inputs, neural operators learn the underlying solution operator that maps input functions (e.g., material property fields, boundary conditions, loadings) to output solution fields. This operator-based learning enables neural operators to generalize across varying initial and boundary conditions without retraining. Additionally, neural operators do not require explicit knowledge of the governing equations, making them particularly well-suited to complex physical phenomena in which a closed-form PDE is not defined. Compared to conventional numerical methods or PINNs, which need to solve the underlying PDEs for each new set of conditions (koric2024deeponet_plasticity; jiang2024mionet_carbon; faroughi2022sciml), neural operators offer significantly more efficient inference based on learned function mapping. These advantages make them particularly well-suited to serve as surrogate models in engineering applications with complex operating conditions.
@@
Among various neural operator architectures, Deep Operator Networks (DeepONet), proposed by lu2021learning, have gained substantial attention due to their formulation. DeepONet is based on the universal approximation theorem for operators (chen1995universal), which establishes its ability to approximate nonlinear solution operators of PDEs given sufficient network complexity and data coverage. This framework allows DeepONet to efficiently learn generalizable operators (lu2022comprehensive).
@@
DeepONet can achieve greater data efficiency, faster training convergence, and improved accuracy (lu2021learning) compared to conventional architectures, such as feed-forward networks (FNNs), sequence-to-sequence models (seq2seq), and residual networks (ResNets) in operator learning tasks. Furthermore, the architecture of DeepONet offers interpretability through its explicit separation of functional roles: the branch net learns to encode the input function, representing the coefficients of a latent basis, while the trunk net learns to output corresponding basis functions over the spatio-temporal domain, which are then combined to form a solution. This design mirrors the structure of functional mappings in infinite-dimensional spaces, where solution operators can be viewed as weighted combinations of basis functions parameterized by input functions (naylor1982linear_operator_theory). This formulation aligns with the mathematical framework of operator learning (kovachki2023neural_operators).
@@
DeepONet has been applied in various fields, such as fluid mechanics (kumar2024deeponet_engine; bai2024deeponet_cfd), material science (yin2022deeponet_solid), structural mechanics (he2023deeponet_structure; ahmed2024deeponet_structure), and thermodynamics (koric2023deepoent_heat; kushwaha2024deeponet_heat). These efforts have highlighted DeepONet’s potential in modeling complex physical processes. However, the potential of DeepONet in geotechnical engineering remains largely unexplored. This study aims to introduce DeepONet as a surrogate modeling approach for geotechnical problems, focusing on the consolidation phenomenon, one of the most prevalent geotechnical engineering problems. The motivation for exploring the consolidation problem is that it can serve as a benchmark to demonstrate DeepONet’s capabilities to the geotechnical community, while also highlighting the role of the different architectural adaptations explored in this study.
@@
More specifically, this study presents a systematic investigation into DeepONet architectural choices tailored for the consolidation phenomenon. We first explain the standard DeepONet architecture, then explore three learning architectures to gain initial insights. The first two models follow DeepONet variants previously investigated (tan2022enhanced_deeponet; jin2022mionet), and the third model adopts a non-traditional architectural choice, with its improved accuracy supported by the analytical solution structure of the consolidation problem. Insights and challenges identified across the explored architectures are shared and used to propose a DeepONet architecture enhanced with TrunkNet Fourier feature embedding, which demonstrates superior performance. We further test the scalability in 3D consolidation scenarios, and present a conceptual application to uncertainty quantification. Overall, this study contributes to the integration of advanced scientific machine learning into geotechnical applications.
@@ 2 Deep operator networks (DeepONet)
DeepONet’s learned basis and corresponding coefficients eliminate the need for manually designed basis functions, which are typically required in traditional numerical methods. For example, the finite element method uses local basis functions defined over a mesh, while spectral methods rely on global functions like sine and cosine series (e.g., Fourier series) chosen in advance (williams2024deeponet_bases). This flexibility allows DeepONet to effectively capture complex, nonlinear operator mappings in high-dimensional PDEs.
@@ 3 Consolidation and associated PDEs
$u(z,t)=\sum_{a=0}^{\infty}\left[\frac{2u_{0}}{A}\sin\left(\frac{Az}{H_{\mathrm{dr}}}\right)\right]e^{-A^{2}T_{v}}$ (5)
Where $A=\frac{\pi}{2}(2a+1)$ under single drainage boundary condition, $u_{0}$ is the initial excess pore pressure, $H_{dr}$ is the drainage path length, and $T_{v}=\frac{C_{v}t}{H_{dr}^{2}}$ is the dimensionless time factor. This analytical form reveals that the solution is composed of spatial sine functions modulated by exponential decay terms dependent on $C_{v}$ and time $t$. This insight informs one of the DeepONet architecture choices that is later explored.
@@ 4 Data generation and training
We use the finite difference method to numerically solve Equation 4. The solution obtained is used to train the DeepONet and to test the performance. The finite difference method first discretizes the spatial domain uniformly into $N_{z}$ nodes separated by a spacing $\Delta z$. It then approximates the spatial second derivative at each internal node $i$ ($1\leq i\leq N_{z}-2$), using a second-order central finite-difference scheme (Equation 6).
@@
When generating data, we consider both uniform and spatially varying initial excess PWP profiles with depth $u(z,0)$ as the initial condition. For the uniform profile, we randomly sample constant functions between 10 kPa and 20 kPa. For the spatially varying (non-uniform) profiles, we sample functions from a Gaussian Random Field (GRF), denoted as $u(z,0)\sim G(\mu_{r},C(z_{1},z_{2}))$. The spatial correlation of this GRF is defined by a Gaussian covariance function:
[…]
Here, $\mu_{r}$ is mean, $C(z_{1},z_{2})$ represents the covariance between two points at depths $z_{1}$ and $z_{2}$. The term $z_{1}-z_{2}||$ is the Euclidean distance between these points. The parameter $l$ is the correlation length (or length-scale), which controls the smoothness of the sampled functions; a larger $l$ generates smoother profiles. The term $\sigma_{r}$ is the standard deviation of the field, controlling the overall magnitude of the covariance. We use a $\sigma_{r}^{2}$ value of $1\,\text{kPa}^{2}$, a correlation length of $0.5$, and sample $\mu_{r}$ uniformly from the range $10$-$20\,\text{kPa}$. The coefficient of consolidation, $C_{v}$, is also sampled uniformly from the range $0.3$-$1.0\,\text{m}^{2}/\text{year}$. The spatial domain spans the normalized depth $\frac{z}{H_{dr}}\in[0,1]$ and normalized time $T_{v}=\frac{C_{v}t}{H_{dr}^{2}}\in[0,2]$.
@@
We train and validate DeepONet using N=40,000 and N=5,000 functions and coefficients, respectively. The trained model is tested on 500 unseen functions and coefficients, with no overlap among the three sets. Specifically, each input function represents an initial excess PWP and is discretized at m=100 fixed, equally spaced sensor locations, denoted by $\{x_{\ell}\}_{\ell=1}^{100}$ (see the left side figure in Figure 2a). That is, for each input function $u^{(i)}$, the branch network receives the vector $\left[u^{(i)}(x_{1}),u^{(i)}(x_{2}),\dots,u^{(i)}(x_{100})\right]$ as input. The corresponding output functions (i.e., solutions $s=\mathcal{G}(u)$) are sampled at P=100 randomly selected evaluation points $y=(z,t)$ (see the red markers in the right side figure in Figure 2a), sampled independently for each input function. These locations are denoted $\{y_{j}^{(i)}\}_{j=1}^{100}$, resulting in a total of $40,000\times 100$ training examples of the form $\left(u^{(i)},y_{j}^{(i)},\mathcal{G}(u^{(i)})(y_{j}^{(i)})\right)$. The validation set is constructed in the same manner, with P=100 output evaluations per input, resulting in $5,000\times 100$ training examples. For testing, we assess the generalization capability of the learned operator $\mathcal{G}_{\Theta}$ on 500 unseen input functions during training. Each is evaluated over a dense $100\times 100$ uniformly spaced grid in the output domain (see the right side figure in Figure 2b), yielding P=10,000 output locations per function to visualize full solution fields. All input and output data are standardized using the mean and standard deviation computed from the training set.
@@ 5 Model architectures
    Model 1 (Figure 3): This model serves as our baseline and adheres to the standard DeepONet formulation introduced by lu2021learning. The branch net processes the input function, which in this case consists of the discretized initial PWP profile, $u(z,0)$. The scalar coefficient of consolidation, $C_{v}$, is concatenated with the discretized $u(z,0)$ vector, forming a single input tensor for the branch net. This design philosophy treats $C_{v}$ as a global conditioning parameter. The implicit assumption is that the branch net should learn a combined representation of how the initial PWP distribution and the material property $C_{v}$ jointly influence the consolidation process. The trunk net takes the spatiotemporal coordinates $(z,t)$ as input.
@@
    Model 2 (Multiple-Input Operator Network) (Figure 4): This architecture modifies the branch net to handle $u(z,0)$ and $C_{v}$ separately before integrating their influence. One branch net is dedicated to encoding the initial PWP function, $u(z,0)$, while a second, auxiliary branch net processes the coefficient of consolidation, $C_{v}$, generating distinct latent representations for each. These two latent representations are then concatenated and passed through a ”merge net” (another small neural network) to produce the final branch output vector. This approach allows the model to potentially learn more disentangled features for the initial condition and the PDE coefficient before their combined effect on the solution basis is determined. jiang2024mionet_carbon reported a performance improvement by introducing the merge net. The trunk net architecture remains the same, taking $(z,t)$ as input. This model architecture aligns with the multiple-input operator network (MIONet) proposed by jin2022mionet.
@@
The performance of Models 1, 2, and 3 is discussed in Section 6, highlighting the challenges to predict excess PWP distribution during the early stages of consolidation, especially for non-uniform distributions across depth. This motivated the authors to explore an additional option. Specifically, an additional architecture option considers Model 3 enhanced with Fourier Feature Embedding (tancik2020fourier), which can help to learn rapidly varying functions (e.g., wang2021eigenvector_ffe, liu2025leveraging).
@@
$\gamma(\textbf{c})=\begin{bmatrix}\sin{(2\pi\mathbf{B}\textbf{c}})\\ \cos{(2\pi\mathbf{B}\textbf{c}})\end{bmatrix}^{T}$ (11)
where c is the input feature vector (trunk net input in this problem) of length $k_{f}$, and $\gamma(\textbf{c})$ is the embedded Fourier feature vector with a length of $2m_{f}$. $\mathbf{B}$ is a $m_{f}$ by $k_{f}$ matrix of frequencies, sampled from $\mathcal{N}(\mathbf{0},\sigma_{f}^{2}I)$. The tunable hyperparameters $m_{f}$ and $\sigma_{f}$ are pre-specified before the neural network training. This embedding introduces periodic features that enrich the input representation, allowing neural networks to approximate functions with high variation. Leveraging the Fourier feature embedding, we propose the following architecture:
@@
All the networks (i.e., branch, trunk, and merge net) in DeepONet models in this study employ fully connected multilayer perceptrons (MLPs). Each MLP comprises 6 layers with 30 neurons per layer. The output dimension of the branch and trunk net for the dot product is set to $q=50$ ( Equation 3). For Fourier feature embedding hyperparameters (Equation 11) in Model 4, we performed a randomized search to select $m_{f}=50$ and $\sigma_{f}=1.0$. Previous studies (e.g., tancik2020fourier; liu2025leveraging) have conducted ablation studies to evaluate the impact of the Fourier feature parameters on the performance of neural networks. Specifically, increasing the value of $m_{f}$ allows neural networks to capture a broader range of frequency components, though this comes at the cost of more computational time. The parameter $\sigma_{f}$ controls the frequency range covered by the Fourier features in the neural tangent kernel. A larger $\sigma_{f}$ allows for the inclusion of higher frequency components, which helps the neural network learn finer details. Conversely, a smaller $\sigma_{f}$ restricts the frequency range, resulting in overly smooth representations that may miss important details. All the DeepONet models are trained using the AdamW optimizer with an initial learning rate of 0.01 and a mini-batch size of 256.